

********************************************************************
UARS WEEKLY STATUS REPORT - 06 NOV 1998
********************************************************************
***********
SOLSTICE
***********
SOLSTICE continues to operate normally, collecting solar data on all
available orbital tracking periods and stellar calibration observations,
whenever possible.
***********
SUSIM
***********
SUSIM successfully gathered its normal solar, offset, and occulation
data during this past week without apparent problems.
Reprocessing of all SUSIM experiment days to version 20 continues.
***********
PEM
***********
Nothing received.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
Nothing received.
***********
MPG
***********
Routine processing of instrument Daily Activity Plans continues
with no problems. The MPG continues to assist the Project
Scientist with the scheduling of instrument operations.
***********
CDHF
***********
The CDHF has received a total of 433 CD-ROM requests as of 06-November-1998.
Three orders were received and shipped.
The CDHF processed current day data through level 3 for all active
instruments. Additionally, CDHF personnel reprocessed: (R)= Requested
(C) = Complete, (F)= Failed (L) = Days remaining to reprocess.
14 days HAL (4.80) PROD (L1) (1873R-1865C-4F=4L)
0 days HAL (4.80) PROD (L3) (1873R-1139C=734L)
49 days SOL (4.62) PROD (L2-L3) (2349R-1963C=386L)
72 days SUS (7.00) PROD (L1-L3) (1249R-72C=1177L)
1 days CLA (7.00) ENG (L1-L3) (18R-4C-2F=12L)
The CDHF installed Y2K compliant CMS v3.9 (DEC/Code Management System)
on CDHF5. Users were able to connect to the existing library and
successfully execute CMS macro commands.
CDHF pending software upgrades: IDL 5.1, OpenIngres, UCSS 10.1, MSS 4.1,
NRT 2.5, UOAS 5.4.
***********
FOT
***********
REPORTING PERIOD
This report covers 23 October 1998 (Orbit 38894, GMT Day 297) through 30
October 1998 (Orbit 38998, GMT Day 303).
SPACECRAFT OPERATION
The observatory is now in Forward Flight (Southern Hemisphere viewing) and
is performing nominally under two-battery operations. The instrument
operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
HALOE 303/2237 38998 Off
The beta angle ranged between 38.1 degrees to 13.9 degrees for this report
period. The beta angle is now increasing toward a minimum angle of 3.0
degrees on 02 November 1998 (DOY 306).
Flight Operations continues to work with the PEM instrument to fine tune
the operational command sequences associated with the OBC Power Monitor
day-only operational capability.
Spacecraft battery 1, 2, and 3 performance monitoring continues. Battery 1
remains OFF the charge relay. Battery 2 and Battery 3 are currently
maintaining greater than 23.8 V end-of-night (EON) load bus voltage.
Battery temperatures are stable with a temperature delta between Battery 2
and Battery 3 of 0.9 to 1.3 degrees C. Additional battery performance data
for each battery on 23 October 1998 and 30 October 1998 are provided in
Appendix D.
The clock error ranged from 10.3 to -11.4 msec during this report period.
Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT for
this report. A summary is listed below and the details of ALERT and Out-
Of-Limits occurrences are listed in Appendix B.
% ALERTS = 00
% OUT-OF-LIMITS = 02
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
% FOT 03
% PEM 07
% HALOE 01
% HRDI 01
% SOLSTICE 01
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
% FOT 06
Two (2) Anomaly Report remained OPEN at the end of this report period (see
Appendix E).
TRANSPONDER FREQUENCIES:
Transponder A: 98/210 - 2287.496941
Transponder B: 98/211 - 2287.499161
GROUND SYSTEM OPERATION
One (1) operational support problems occurred this period. Anomaly Reports
are included in Appendix E.
The data loss calculations supplied by Data Capture Facility (DCF) from
mission start through UTC Day 98/302 show the observatory data loss to be
21 hours, 11 minutes, 06.740 seconds (an increase of 31.68 seconds since
last report period). This is a 0.0340 percent data loss which equals a
99.9660 percent data capture for the mission.
The increase of 31.68 seconds occurred over seven (7) days:
7 3.136 seconds on Day 296 due to hits in data and insufficient overlap;
7 2.368 seconds on Day 297 due to hits in data and line errors (IP);
7 0.632 seconds on Day 298 due to hits in data and IP hits;
7 14.368 seconds on Day 299 due to line errors (IP) hits in data and
insufficient overlap;
7 3.200 seconds on Day 300 due to line errors (IP) and data hits;
7 6.240 seconds on Day 301 due to hits in data; and
7 0.736 seconds on Day 302 due to hits in data (IP).
The FOT modified the CMS resident Relative Time Sequence 06 (RTS 06) to
facilitate proper PMON-controlled day-only operation of the PEM instrument.
A request from SUSIM Operations to perform another series of SSPP Gridded
Offset maneuvers on GMT days 311 through 316 was received this week. The
offsets will utilize the 5 and 10 arcminute step intervals, as opposed to
the 15 and 30 arcminute intervals
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
FOT continues to improve a power profile for UARS based on the current
status of the Modular Power System.
Investigation continues into possible impacts of the Leonid meteorite
shower in November 1998.
Investigating the feasibility of a new PMON action to change battery V/T
curves based on loading to mitigate battery risks in sun pointing safehold.
YEAR 2000 OPERABILITY SUPPORT
UARS Test and Training Simulator (UTTS) - No change.
User Planning System (UPS) - No change.
Multi Satellite Operations Control Center (MSOCC) - No change.
Command Management System (CMS) - The FOT received a valid Y2K Star Catalog
from FDF this week. Apparently, the star catalog files that FDF was
sending to the FOT previously were being incorrectly re-formatted by the
FORMATS system. Eventually, FDF did the necessary file re-formatting in
house and bypassed FORMATS altogether by sending the re-formatted star
catalog file directly to the CMS. Once the file was received on CMS, it
was ran through the CMS poller, which performed all of the necessary
conversions to the CMS readable Raw Data Base RDB file.
Still waiting for a 8-week S/C Orbital Events file from FDF which contains
a SC_YAWMAN event. Once the FOT receives this file, the last remaining Y2K
test function (#12) of the 1999-to-2000 test case (#1) can be completed.
In addition, the FOT is also awaiting a valid set of 3-day PSAT and S/C
Orbital Events files from FDF which span the 2000-to-2001 century
crossover. Up to this point, all of the PSAT and S/C Orbital Events files
generated for that time frame appear to have an Rascending nodeS
inconsistency between the two files, making it impossible to process them
on CMS.
Flight Dynamics Facility (FDF) - see CMS paragraph.
IP NETWORK TESTING
Space Network (SN) - Operational.
NCC 98 system - In planning.
Deep Space Network (DSN) and Ground Spaceflight Tracking and Data Network
(GSTDN) - No change. Coordinating test with NIA personnel.
UARS PROPULSION MODULE B/D THRUSTER LINE FILL
GSFC Propulsion Branch reported that the propulsion engineers will be ready
to support the UARS simulated line refill test after 04 November 1998. FOT
is developing the command techniques to achieve a 16 ms and an 8 ms command
timing of the latch valve between on and off. This will be checked with
the simulator before using the latch valve test hardware. Jackson and Tull
were funded to built the test hardware but do not have funds to support
operation initial checkout or operation of the test.
UARS RE-ENGINEERING - TELEMETRY PROCESSING
Flight operations is providing assistance on both sustaining engineering
and re-engineering activities to incorporate needed features into the
Trending system. A meeting was held on 30 October 1998 with the sustaining
engineering team regarding DR and ER items. Supporting details regarding
current trending products and desired functionality of the new trending
system were provided.
MEETINGS
A meeting was held 29 October 1998 with Mike Bay (Jackson and Tall),
Dimitrios Mantziaras, and Dan Wise to discuss the best way to command the
Latch valves for the UARS line fill test. Information was received on the
RIU modes which will be helpful to the test. An agreement as to the
technique (PDB and ATC commands) was reached.
The November battery telecon is scheduled for 24 November 1998. The 29
October 1998 telecon was not attended by the UARS FOT due to conflicts.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw Around Maneuver is scheduled for scheduled for 02
November 1998.
UARS will pass through the Leonid meteor shower on 17 November 1998 at
19:43 UTC (+/- 3 hours).
The United States Naval Observatory (USNO) issued a Time Service
announcement advising that the International Earth Rotation Service (IERS)
will retard UTC time by 1.0 second at the end of December 1998. This means
the addition of one leap second as the last second of 1998.
The next annular solar eclipse will occur on 16 February 1999 over the
Indian Ocean, Australia and the South Pacific Ocean.
PERSONNEL CHANGES
None
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 30 October 1998
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ -------
38903 297/1518 98-408 PEM PEM Turn-on test
38918 298/1522 98-409 OBC Clk Adj fr 0.00 to 20.6 msec/day
38937 299/2040 98-410 PEM PEM Turn-On test
38937 299/2141 98-411 PEM PEM Turn-On test
38942 300/0607 98-412 OBC Clk Adj fr 20.6 to 0.00 msec/day
38951 300/2030 98-413 OBC PEM/HEPS test
38952 300/2145 98-414 SOLS Reset Observation List
38967 301/2046 98-415 PEM PEM/HEPS
38982 302/2052 98-416 PEM PEM/HEPS Test
38994 303/1734 98-417 PEM RTS-6 Activate
38997 303/2157 98-418 PEM PEM/HEPS Test
38997 303/2157 98-419 HRDI HRDI Microload
38998 303/2337 98-420 HAL0 Power OFF
_______________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 30 October 1998
MNEMONIC STATE DESCRIPTION ORBITS
--------- ----- ------------ ------
NBRECBI Y-HI NBTR B Total Current 38900
HRPTSPBIMON Y-HI PTS Pulse Bus Current Monitor 38950
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
--------- ------------ ------
None this report period
_______________________________________________________________
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 30 October 1998
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ---- -----------------
38918 298/15:26 (25 Oct 98) from 00.0 to 20.6 msec/day
38942 300/06:10 (27 Oct 98) from 20.6 to 00.0 msec/day
_______________________________________________________________
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 30 October 1998
23 Oct 1998 - GMT Day 296 Beta = 38.1 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------- ------
1 +728.0/+728.0 -3.05/-3.99 +0.0/ +0.0 17.9 0.000 0.0
2 +44.8/-5.6 +5.47/+4.51 +29.2/-14.4 26.1 1.028 20.6
3 +134.4/0.0 +4.51/+3.23 +28.4/-14.4 26.1 1.026 20.2
30 Oct 1998 - GMT Day 303 Beta = 13.9 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------- ------
1 +470.0/+190.4 -3.05/-3.99 +0.0/ +0.0 17.1 0.000 0.0
2 +44.8/ 0.0 +7.09/+5.47 +31.6/-14.4 24.2 1.022 22.8
3 +128.8/+11.2 +5.79/+4.51 +30.8/-16.0 24.2 1.020 22.6
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- -------- ----------
no changes (V/T 5)
APPENDIX E
UARS ANOMALY REPORTS
Week ending 30 October 1998
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
98-083 FOT 38937 299/2040 Playback Errors
98-084 FOT 38953 300/2327 TAC 8 Crashed
98-085 FOT 38971 302/0431 Multiple TAC crashes
98-086 FOT 38978 302/1457 TAC 2 Crashed
98-087 FOT 38994 303/1637 TAC #2 Crashed
98-066 FOT 38998 303/2337 14170 Generic Late Acquisition
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
98-082 FOT 38888 296/15:14 20288 IP TELEMETRY HITS
98-083 FOT 38937 299/20:40 NBR P/B ERRORS
98-084 FOT 38953 300/23:27 TAC 8 CRASH
98-085 FOT 38971 302/04:31 TAC 8 CRASHES
98-086 FOT 38978 302/14:57 TAC 2 CRASH
ANOMALY INVESTIGATION REPORTS GENERATED
AIR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------- ------ ----- ---- --- -----------------
None this report period
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
98-066 FOT Oct 98 See List 14170 GENERIC LATE ACQ
98-087 FOT` 38994 303/16:37 TAC 2 CRASH
SEVNTFW44